OpenStudy (anonymous):

Find dy/dx by implicit differentiation. (sin πx + cos πy)^9 = 34

OpenStudy (anonymous):

could have sworn we did this already. no?

OpenStudy (anonymous):

hey sorry my internet is acting up

OpenStudy (anonymous):

yeah but we never finsihed it

OpenStudy (anonymous):

oh the last job was just algebra, so solve that equation for \[y'\] same thing we still have to do in other post

OpenStudy (anonymous):

i dnt knwo where that post has gone :(

OpenStudy (anonymous):

satellite i know this sounds needy but do you just mind giving me the answer for this problem and the other one...my homework is timed and i only have 10mins left:(please

OpenStudy (anonymous):

hold on it will take me a minute or 5 to do the algebra

OpenStudy (anonymous):

i feelso bad but im desperate

OpenStudy (anonymous):

no problem let me do it

OpenStudy (anonymous):

oki holding on thanks:)

OpenStudy (anonymous):

ok you here. answer not to this one but to the last one, after a bunch of algebra i got \[y''=\frac{48y^5x^7}{30x^8y^4+4y^{10}}\]

OpenStudy (anonymous):

okiay leme punch that in

OpenStudy (anonymous):

god i hope i did the algebra right because i did not check. is this a test?

OpenStudy (anonymous):

anyway this one is easier. this one is \[y'=\frac{\cos(\pi x)}{\sin(\pi y)}\]

OpenStudy (anonymous):

its graded homework

OpenStudy (anonymous):

i can check the algebra again if you like

OpenStudy (anonymous):

it is long

OpenStudy (anonymous):

yeah if you dont mind

OpenStudy (anonymous):

ok in the mean time i think the other is right. hold on

OpenStudy (anonymous):

the other one was right!!thankyou so much:)

OpenStudy (anonymous):

christ this algebra is a pain. try this for the last one

OpenStudy (anonymous):

sorry yeah its a real pain

OpenStudy (anonymous):

numerator \[48y^6x^7-45x^{16}\] denominator \[4y^{11}\]

OpenStudy (anonymous):

no guarantees

OpenStudy (anonymous):

its k leme try

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

it was right!ur a genious..omg thank you so much!!!

OpenStudy (anonymous):

it was right?!

OpenStudy (anonymous):

wow. invoice is in the mail